OpenStudy (anonymous):

find the limits when h ->0 h /tan^2h

OpenStudy (anonymous):

h approaches 0 and tan^2h approaches 0 as h approaches 0, so you have an indeterminate form. Use L'Hopital's rule to obtain\[\lim_{h \rightarrow 0^+}\frac{h}{\tan^2h}=\frac{\lim_{h \rightarrow 0^+}h}{\lim_{h \rightarrow 0^+}\tan^2h}=\frac{\lim_{h \rightarrow 0^+}1}{\lim_{h \rightarrow 0^+}2\tan (h) \sec^2 h}=\frac{1}{\lim_{h \rightarrow 0^+}2\tan (h) \sec^2 h}\rightarrow +\infty\]

OpenStudy (anonymous):

Refresh the screen if it doesn't show up properly.

OpenStudy (anonymous):

where did u get sec

OpenStudy (anonymous):

That's the limit from the right (i.e. heading toward 0 from the positive part of the number line).

OpenStudy (anonymous):

From the derivative of tan(x):\[\frac{d}{dx}\tan^2x=2\tan x \sec ^2 x\](chain rule)

OpenStudy (anonymous):

i didnt know u take the derivative

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0^-}\frac{h}{\tan^2h}=\frac{\lim_{h \rightarrow 0^-}h}{\lim_{h \rightarrow 0^-}\tan^2h}=\frac{\lim_{h \rightarrow 0^-}1}{\lim_{h \rightarrow 0^-}2\tan (h) \sec^2 h}=\frac{1}{\lim_{h \rightarrow 0^-}2\tan (h) \sec^2 h}\rightarrow -\infty\]

OpenStudy (anonymous):

You take the derivative for L'Hopital's rule.

OpenStudy (anonymous):

In its original form, you can't say anything about the limit. Anything where you have\[\frac{0}{0}, \pm \frac{\infty}{\infty}\]is an indeterminate form (you can't say anything about it). We use (usually) L'Hopital's rule to reduce it into a determinable form.

OpenStudy (anonymous):

Here: http://en.wikipedia.org/wiki/L'H%C3%B4pital's_rule

OpenStudy (anonymous):

ohhhh,, nice

OpenStudy (anonymous):

so when u do that?

OpenStudy (anonymous):

when u cant do anything about the limit

OpenStudy (anonymous):

If you look above, I tell you when you have an indeterminate form.

OpenStudy (anonymous):

0/0 +/ infty/infty

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Here, the h in the numerator approached 0 the tan^2h in the denominator approached 0 when h approached 0, so you had a case of 0/0 which is indeterminate.

OpenStudy (anonymous):

You look at the numerator and denominator separately first (i.e. to see if you get an indeterminable form). If it's the case, apply L'Hopital's rule or something else (but no one uses the 'something else').

OpenStudy (anonymous):

jajjaaj

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks.. very good explanation

OpenStudy (anonymous):

welcome